OpenStudy (anonymous):

The measure of two complementary angles is in the ratio 1 : 9. What are the degree measures of the two angles?

OpenStudy (anonymous):

What do complementary angles mean?

OpenStudy (anonymous):

they compliment each other

OpenStudy (anonymous):

ah? so one angle says to the other " hey, nice angle" ?

OpenStudy (anonymous):

No they both equal to 90 degrees

OpenStudy (anonymous):

and the other says, "Thanks, you look quite good yourself" ? Is that what you mean by they compliment each other?

OpenStudy (anonymous):

okay good.

OpenStudy (anonymous):

that means they add up to 90 degrees, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right. so let us say one angle is A and the other angle is B

OpenStudy (anonymous):

A : B = 1 : 9, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and A+B = 90. Are you with me?

OpenStudy (anonymous):

ya i get you so far

OpenStudy (anonymous):

good. so you have two equations A + B = 90 and A/B = 1/9

OpenStudy (anonymous):

A / B = 1 / 9 => A = 9B

OpenStudy (anonymous):

substitute A in the other equation and solve for B